# Chinese Academy of Sciences, Beijing 100190, China

### v1

@@ 6 Combined Analysis of Primordial Neutrino Asymmetries
In previous sections, we have shown that the primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving significant observable imprints on the Universe, such as the abundance of light elements produced during the BBN, the CMB anisotropies, and the BAO imprints in the LSS. These cosmological observations can then be used to constrain the primordial neutrino asymmetries. In this section, we consolidate the results of the previous sections and discuss the constraints on the primordial neutrino asymmetry derived from contemporary observations of the BBN, CMB and BAO. In addition, we briefly explore the corresponding constraints on UV model parameters using a specific model that generates primordial neutrino asymmetries in the early universe.
@@ 6.1 Observational Constraints on Primordial Neutrino Asymmetry
Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon abundance $\omega_{b}$, the cold dark matter abundance $\omega_{cdm}$ and the total neutrino mass $\sum m_{\nu}$ from various datasets and cosmological models.
In this work we utilize the following datasets for BBN, CMB, and BAO in our analysis:
BBN
The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance adopted by the PDG group, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , which is based on the results of Cooke:2017cwo .
CMB
@@
We note that the helium abundance from the BBN data set primarily constrains the primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$. The BAO dataset is mainly used to constrain the total neutrino mass $\sum m_{\nu}$, but as we will discuss, it also places additional constraints on primordial neutrino asymmetries.
@@
To extract the constraints from these datasets, we conduct MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx . For analysis, the $N_{\rm eff}$ is adopted as Eq. (63), derived in Section 3. The helium and deuterium abundances predicted by BBN are computed using our fitting formulas in Equation (85). As shown in Fig. 10, the effects of neutrino spectral distortions on the CMB angular power spectrum and the BAO in the matter power spectrum are small. Therefore, for simplicity, the (anti)neutrino spectral distortions are neglected in the MCMC analysis. The parameter set for the CMB and BAO datasets includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$, where $\theta_{s}$ represents the sound horizon, $A_{s}$ is the amplitude of the primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization. On the other hand, we only consider $\{\omega_{b},\xi_{\nu}\}$ for the BBN dataset, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
@@
Figure 11: 2D and 1D marginalised posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$, and scalar spectral index $n_{s}$. The red, blue, and green regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
@@
Figure 12: 1D marginalised distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
@@
With the aforementioned choices of parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis. The Gelman-Rubin criterion for all chains converged to $R-1<0.01$. The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mas are also presented from the constraints on the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$. For the same model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig. 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
@@
Concerning the constraints on the primordial neutrino asymmetries, we see that the recent EMPRESS data in the BBN dataset points to a positive primordial neutrino asymmetry $\xi_{\nu}=0.039\pm 0.013$, which is different from zero with a $3.0\,\sigma$ significance. Combining the BBN and CMB data, the resulting constraints on $\xi_{\nu}$ are significantly reduced to $\xi_{\nu}=0.028\pm 0.012$, different from zero with a $2.3\,\sigma$ significance. Compared to the constraints $\xi_{\nu}=0.034\pm 0.014$ reported in previous study with the same BBN and CMB datasets Escudero:2022okz , our result indicate a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty. Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results, while the significance of a positive neutrino asymmetry is reduced to $2.0\,\sigma$. Our study demonstrate the additional constraining power of the BAO experiments on the primordial neutrino asymmetry, which are not aware in previous studies.
@@
We find that the constraints on other parameters also vary significantly with different datasets, especially for $\omega_{b}$, $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$, as shown in Fig. 11. For the constraints on $\omega_{b}$, the difference is largely due to the fact that we chose the nuclear reaction rates in PRIMAT in the process of BBN, which predicted a deuterium abundance lower than the current constraints from PDG, thus leading to significantly lower constraints on $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While the constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in  Escudero:2022okz , a detailed study with different nuclear reaction rates will be performed in our companion paper. The inclusion of the BOSS BAO data significantly changes the corresponding constraints on $\omega_{cmb}$, $n_{s}$ and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late universe. Especially, the resulting $95\%$ upper limit for the total mass of neutrinos are greatly tightened from $0.414$ eV by BBN + CMB to $0.131$ eV by BBN + CMB + BAO, since the BAO data help to break the CMB degeneracy. Recently, the BAO observations of DESI have put an even stronger bound on neutrino masses $\sum m_{\nu_{i}}\lesssim 0.07\,eV$ DESI:2024mwx . The relevant studies on primordial neutrino asymmetry and neutrino masses with DESI BAO data will also be discussed in our companion paper.
@@
Finally we briefly discuss the effects of different parameter sets on the resulting constraints, especially the interplay between the primordial neutrino asymmetry and the total mass of neutrinos. Using the data set of BBN + CMB + BAO, we present in Fig. 12 the resulting 1d marginalised distributions for three choices of the parameter set, first is the set of $\Lambda CDM+m_{\nu_{1}}$, second is the $\Lambda CDM+\xi_{\nu}$, while the third is $\Lambda CDM+m_{\nu_{1}}+\xi_{\nu}$. As we can see, while the different choices of the parameter set lead to quite modifications to the constraints on the six parameters of $\Lambda CDM$, especially for $\omega_{b}$ and $n_{s}$, the resulting constraints on the additional parameters are almost the same for different parameter sets. The constraints on the total mass of neutrinos are only slightly tightened while including primordial neutrino asymmetry $\xi_{\nu}$ as a free parameter, while the constraints on $\xi_{\nu}$ are almost the mass as the inclusion of free $\sum m_{\nu}$. Therefore, we conclude that the constraints on neutrino masses and the primordial neutrino asymmetries can indeed be analysed separately if one do not concern the constraints on the other parameters in $\Lambda CDM$, otherwise the parameter sets with both the neutrino masses and the primordial neutrino asymmetries must be considered.
@@ 6.2 Constraints on UV model parameters
With the constraints on the primordial neutrino asymmetries, we discuss in this subsection the corresponding constraints on the UV model parameters with a specific model to generate the primordial neutrino asymmetries. Here we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced in the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which utilizes flat directions in the minimal supersymmetric standard model (MSSM). When some flat direction with lepton number has a large field value during inflation, it generates lepton number by field dynamics after inflation. At the same time, Q-balls are produced and the generated lepton number is all confined inside the Q-balls. Thus, such Q-balls have a large lepton number and are called L-balls. Since the sphaleron effect is inactive inside the L-balls, the lepton asymmetry produced by the Affleck-Dine mechanism are kept without being converted to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg . The L-balls decay into neutrinos below the electroweak scale and realize a large lepton asymmetry in the universe.
Following the results of Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio between the difference of the number density of neutrinos over antineutrinos and the photon number density,
@@
where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density and $g_{\star}=10.75$ is the relativistic degree of freedom at the neutrino decoupling epoch. In addition, the resulting $\eta_{L}$ in the Q-ball decay model can be expressed as
@@
where $m_{3/2}$ is the the gravitino mass, $T_{\mathrm{D}}$ is the decay temperature of the L-balls and can be expressed with the other UV model parameters with
@@
where $g_{*}(T_{\mathrm{D}})$ is the relativistic number of degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, $N_{\ell}$ is the number of decay channels. Therefore, once we have obtained the constraints on the lepton asymmetry from cosmological observations, they can be directly translated into constraints on the model parameters of the UV model.
According to our study, the constraints on $\eta_{L}$ from the BBN + CMB + BAO datasets are $\eta_{L}=0.0025\pm 0.0013$. Therefore, the constraint on the L-balls decay temperature is
@@
These constraints can also be converted into the constraints on other UV model parameters, such as the SUSY breaking scale, as seen in the preliminary study of Kawasaki:2022hvx .
@@ 7 Conclusion
The presence of primordial neutrino asymmetries before the neutrino decoupling epoch has significant implications for both the neutrino decoupling process and subsequent cosmological phenomena, and can therefore be constrained by observations of BBN, CMB and LSS. However, the effects of primordial neutrino asymmetries on neutrino decoupling—particularly concerning the resulting $N_{\rm eff}$ and the spectral distortions of (anti)neutrinos, including the impacts of flavor oscillations and FTQED corrections—remain ambiguous. These ambiguities affect the interpretation of subsequent processes. Consequently, studies of primordial neutrino asymmetries in relation to BBN and the CMB often employ approximations, such as using thermal Fermi-Dirac distributions for (anti)neutrinos, without thorough scrutiny.
@@
In this work, we aim to address these ambiguities by numerically solving the neutrino decoupling process and analyzing the implications of the resulting (anti)neutrino density matrices for BBN, CMB, and LSS. We begin by re-deriving the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Focusing on primordial neutrino asymmetries with complete flavor equilibration, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv , including the effects of neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then evaluated based on the resulting density matrices, and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$, while the second and last term represent the corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and from the non-instantaneous effect, respectively. We also confirm that the spectral distortions of neutrinos and antineutrinos with non-zero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with the priliminary study in Grohs:2016cuu , where neutrino spectral distortions were evaluated while neglecting flavor oscillations and FTQED corrections.
@@
Using the density matrices for neutrinos and antineutrinos, we further investigate the implications of primordial neutrino asymmetries for BBN, CMB and LSS. For BBN, we introduce a novel correction to the weak interaction rates arising from the distinct spectral distortions of neutrinos and antineutrinos. The light element abundances are then computed with the public BBN code PRIMAT Pitrou:2018cgg , incorporating the other additional corrections on weak rates. According to the the EMPRESS measurement on the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results immediately indicate a positive primordial neutrino asymmetry $0.032\leq\xi_{\nu}\leq 0.052$ with fixed baryon abundance $\omega_{b}$. The effects of the new proposed correction are also examined, and we find that the spectral distortions for neutrinos and antineutrinos are important for precise predictions for BBN, especially for the abundance of Helium-4. Compared with the results in the instantaneous limit, the traditional treatment with Fermi-Dirac distributions predict a negative relative difference for a positive primordial neutrino asymmetry, while our precise treatment with spectral distortions show that the relative difference is still positive. Additionally, we have provided semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be useful for future studies. For the CMB and LSS, a novel finding of our study is that primordial neutrino asymmetries also affect the amplitudes of the acoustic oscillations, both for the CMB and for baryon acoustic oscillations (BAO). This shows that BAO observations could serve as a tool for constraining primordial neutrino asymmetries.
@@
Based on our systematic understanding of primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations. Combining the EMPRESS BBN and Planck CMB datasets, the primordial neutrino asymmetry constraint yields $\xi_{\nu}=0.028\pm 0.013$, indicating a $\sim 18\%$ smaller central value and a slight reduction in the uncertainty compared to previous results $\xi_{\nu}=0.034\pm 0.014$ obtained in Escudero:2022okz . Incorporating BOSS BAO observations in the MCMC analysis, the primordial neutrino asymmetry constraint further reduce to $\xi_{\nu}=0.024\pm 0.013$, indicating a $\sim 29\%$ reduction in the central value compared to previous results and a $2.0\,\sigma$ significance of a positive neutrino asymmetry. We also discuss the impacts of treating the sum of neutrino mass $\sum m_{\nu}$ and the primordial neutrino asymmetry $\xi_{\nu}$ as free parameters on the resulting constraints for each of them. Finally, we also show how the constraints on the neutrino asymmetries can be translated into constraints on the UV model parameters with a specific model in which the large neutrino asymmetries are generated by the decay of Q-balls Kawasaki:2002hq ; Gelmini:2020ekg .
@@
Our work provides the framework to obtain the implications of new physics beyond the SM and $\Lambda$CDM model for neutrino decoupling and subsequent processes, and the corresponding cosmological constraints from present and future BBN, CMB and LSS observations. Thus, our work facilitates a more comprehensive analysis for primordial neutrino asymmetry and other models with new physics, such as precise consideration of the neutrino synchronous oscillations Froustey:2021azz ; Froustey:2024mgf or neutrino non-standard interactions Du:2021idh ; Du:2023upj , based on more and more precise cosmological observations in the future.

### v2

@@ 6 Combined Analysis of Primordial Neutrino Asymmetries
In previous sections, we have demonstrated that primordial neutrino asymmetry significantly influences the neutrino decoupling process, leaving observable imprints on the Universe, such as the abundance of light elements produced during BBN, the anisotropies in the CMB, and the BAO features in the LSS. These cosmological observations can be used to constrain primordial neutrino asymmetries. In this section, we consolidate the results from previous sections and discuss the constraints on primordial neutrino asymmetries derived from present BBN, CMB, and BAO observations. Furthermore, we briefly explore the associated constraints on UV model parameters by using a specific model that generates primordial neutrino asymmetries in the early universe.
@@ 6.1 Observational Constraints on Primordial Neutrino Asymmetry
Table 3: Summary of constraints on the primordial lepton asymmetry $\xi_{\nu}$, the baryon density $\omega_{b}$, the cold dark matter density $\omega_{cdm}$, and the total neutrino mass $\sum m_{\nu}$ in normal hierarchy from various datasets and cosmological models.
The datasets used for BBN, CMB, and BAO analyses in this work are as follows:
BBN:
The EMPRESS measurement of helium abundance, ${Y\rm_{P}}=0.2370^{+0.0034}_{-0.0033}$ empress , and the deuterium abundance from the PDG, $10^{5}\times{\rm D/H}=2.547\pm 0.025$ pdg , based on Cooke:2017cwo .
CMB:
@@
It should be noted that the helium abundance from the BBN dataset primarily constrains primordial neutrino asymmetries, while the deuterium abundance is more sensitive to the baryon density $\omega_{b}$. The BAO dataset primarily constrains the total neutrino mass $\sum m_{\nu}$, but also provides additional constraints on primordial neutrino asymmetries.
@@
To extract constraints from these datasets, we perform MCMC analyses using the Montepython engine Audren:2012wb ; Brinckmann:2018cvx . For the analysis, $N_{\rm eff}$ is derived from Eq. (63) as discussed in Section 3. The predicted helium and deuterium abundances from BBN are calculated using our fitting formulas in Eq. (86). As shown in Fig. 10, the impact of neutrino spectral distortions on the CMB angular power spectrum and BAO in the matter power spectrum is minor. Therefore, for simplicity, (anti)neutrino spectral distortions are neglected in the MCMC analysis. For the CMB and BAO datasets, the parameter set includes the six independent parameters of the $\Lambda$CDM model, along with the neutrino degeneracy parameter $\xi_{\nu}$ and the mass of the lightest neutrino in the normal hierarchy $m_{\nu_{1}}$, i.e., $\{\omega_{b},\omega_{cdm},\theta_{s},A_{s},n_{s},\tau_{\mathrm{reio}},\xi_{\nu},m_{\nu_{1}}\}$. Here, $\theta_{s}$ is the sound horizon, $A_{s}$ is the amplitude of primordial curvature perturbations at the pivot scale $k=0.05\,\mathrm{Mpc}^{-1}$, $n_{s}$ is the scalar spectral index, and $\tau_{\mathrm{reio}}$ is the optical depth to reionization. For the BBN dataset, we consider only $\{\omega_{b},\xi_{\nu}\}$, as the predicted light element abundances from BBN are largely independent of other cosmological parameters within the relevant range.
@@
Figure 11: 2D and 1D marginalized posterior distributions for the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the scalar spectral index $n_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$. The green, blue and red regions represent constraints from BBN, CMB + BAO, and their combination, respectively.
@@
Figure 12: 1D marginalized distributions for different parameter sets derived from the BBN + CMB + BAO datasets.
@@
Using the specified parameters and datasets, we modified the MontePython sampler and CLASS codes to perform the MCMC analysis. All chains converged with a Gelman-Rubin criterion of $R-1<0.01$. The resulting $1\,\sigma$ cosmological constraints are summarized in Table 3, where the $95\%$ confidence limits on the total neutrino mass are also presented, derived from constraints on the mass of the lightest neutrino for the normal hierarchy, $m_{\nu_{1}}$. For the model $\Lambda$CDM + ${\xi_{\nu}}$ + ${m_{\nu_{1}}}$, we present the 1D and 2D marginalized posterior distributions of key parameters that show significant variations across different datasets in Fig. 11, including the baryon abundance $\omega_{b}$, the neutrino degeneracy parameter $\xi_{\nu}$, the cold dark matter abundance $\omega_{cdm}$, the sound horizon $\theta_{s}$, and the mass of the lightest neutrino for the normal hierarchy $m_{\nu_{1}}$.
@@
Regarding constraints on primordial neutrino asymmetries, the EMPRESS data in the BBN dataset indicates a positive primordial neutrino asymmetry, $\xi_{\nu}=0.039\pm 0.013$, which differs from zero at $3.0\,\sigma$ significance. Combining BBN and CMB data significantly reduces the constraint to $\xi_{\nu}=0.028\pm 0.012$, which differs from zero at $2.3\,\sigma$ significance. Compared to previous results ($\xi_{\nu}=0.034\pm 0.014$) using the same BBN and CMB datasets Escudero:2022okz , our results show an 18% reduction in the central value and a slight decrease in uncertainty. Incorporating the BOSS BAO observations further reduces the constraint on primordial neutrino asymmetry to $\xi_{\nu}=0.024\pm 0.012$, representing a 29% reduction in the central value relative to previous findings, while the significance of a positive neutrino asymmetry decreases to $2.0\,\sigma$. Our analysis highlights the additional constraining power of BAO experiments on primordial neutrino asymmetries, which was not fully explored in previous studies.
@@
We find that constraints on other parameters, such as $\omega_{b}$, $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$, also vary significantly depending on the dataset used, as shown in Fig. 11. For $\omega_{b}$, the discrepancy is mainly attributed to the choice of nuclear reaction rates in the PRIMAT code during the BBN process, which predicts a deuterium abundance lower than current constraints from PDG, leading to a significantly lower $\omega_{b}$ Pitrou:2018cgg ; Burns:2022hkq ; Escudero:2022okz . While constraints on the primordial neutrino asymmetry $\xi_{\nu}$ do not strongly depend on the choice of nuclear reaction rates, as discussed in Escudero:2022okz , a detailed study using different nuclear reaction rates will be presented in our companion paper. The inclusion of BOSS BAO data significantly modifies the constraints on $\omega_{cdm}$, $n_{s}$, and $m_{\nu_{1}}$ due to its sensitivity to the evolution of the late Universe. In particular, the $95\%$ upper limit on the total neutrino mass tightens from $0.414$ eV (BBN + CMB) to $0.131$ eV (BBN + CMB + BAO), as BAO data help break the CMB degeneracy. Recently, DESI BAO observations have placed even stronger constraints on neutrino masses, $\sum m_{\nu_{i}}\lesssim 0.07\,\mathrm{eV}$ DESI:2024mwx . Relevant studies on primordial neutrino asymmetry and neutrino masses using DESI BAO data will be addressed in our companion paper.
@@
Finally, we briefly discuss the impact of different parameter sets on the resulting constraints, particularly the interplay between primordial neutrino asymmetry and total neutrino mass. Using the BBN + CMB + BAO dataset, we present in Fig. 12 the 1D marginalized distributions for three parameter sets: $\Lambda$CDM + $m_{\nu_{1}}$, $\Lambda$CDM + $\xi_{\nu}$, and $\Lambda$CDM + $m_{\nu_{1}}$ + $\xi_{\nu}$. As shown, while different parameter sets lead to significant modifications in constraints on the six $\Lambda$CDM parameters, especially $\omega_{b}$ and $n_{s}$, the resulting constraints on additional parameters are consistent across different parameter sets. The corresponding constraints from different parameter sets are also summarized in Table 3. We find that the constraints on the total neutrino mass are only slightly tightened when $\xi_{\nu}$ is included as a free parameter, while constraints on $\xi_{\nu}$ remain nearly unchanged when $\sum m_{\nu}$ is a free parameter. Therefore, we conclude that constraints on neutrino masses and primordial neutrino asymmetries can be analyzed independently if one is not concerned with constraints on other $\Lambda$CDM parameters. However, if constraints on other $\Lambda$CDM parameters are of interest, a parameter set including both of neutrino mass and primordial neutrino asymmetry must be considered.
@@ 6.2 Constraints on UV Model Parameters
In this subsection, we discuss the constraints on UV model parameters based on the constraints on primordial neutrino asymmetries. Specifically, we consider a Q-ball decay model Kawasaki:2022hvx , where Q-balls are produced through the Affleck-Dine mechanism Kusenko:1997si ; Enqvist:1997si ; Kasuya:1999wu , which operates along flat directions in the minimal supersymmetric standard model (MSSM). When a flat direction associated with lepton number reaches a large field value during inflation, it generates a lepton number via field dynamics after inflation. At the same time, Q-balls form, and the generated lepton number is confined within them. These Q-balls, referred to as L-balls, possess a large lepton number. Since the sphaleron process is inactive inside L-balls, the lepton asymmetry generated by the Affleck-Dine mechanism is preserved without conversion to a baryon asymmetry Kawasaki:2002hq ; Gelmini:2020ekg . After the electroweak scale, the L-balls decay into neutrinos, resulting in a significant lepton asymmetry in the universe.
Following Kawasaki:2022hvx , the primordial lepton asymmetry $\eta_{L}$ is defined as the ratio of the difference between the number densities of neutrinos and antineutrinos to the photon number density:
@@
where $s=\frac{2\pi^{2}}{45}g_{\star}T_{\gamma}^{3}$ is the entropy density, and $g_{\star}=10.75$ is the relativistic degree of freedom at neutrino decoupling. In the Q-ball decay model, $\eta_{L}$ can also be expressed as:
@@
where $m_{3/2}$ is the gravitino mass and $T_{\mathrm{D}}$ is the L-balls’ decay temperature, which is a function of other UV model parameters:
@@
where $g_{*}(T_{\mathrm{D}})$ is the relativistic degrees of freedom at $T_{\mathrm{D}}$, $\beta\simeq 6\times 10^{-4}$ is a dimensionless constant, $M_{F}$ is the SUSY breaking scale, and $N_{\ell}$ represents the number of decay channels. Hence, the cosmological constraints on lepton asymmetry can be translated into constraints on the UV model parameters.
Our analysis of the BBN + CMB + BAO datasets yields a constraint on the lepton asymmetry of $\eta_{L}=0.0025\pm 0.0013$. The corresponding constraint on the L-balls’ decay temperature is:
@@
These constraints can be further translated into constraints on other UV model parameters, such as the SUSY breaking scale, as shown in Kawasaki:2022hvx .
@@ 7 Conclusion
The existence of primordial neutrino asymmetries prior to the neutrino decoupling epoch has significant implications for the neutrino decoupling process and subsequent cosmological phenomena. These asymmetries can therefore be constrained through observations of BBN, CMB, and LSS. However, the effects of primordial neutrino asymmetries on neutrino decoupling, particularly with respect to the resulting $N_{\rm eff}$ and spectral distortions of (anti)neutrinos—including the impacts of flavor oscillations and FTQED corrections-remain uncertain. These uncertainties make it difficult to interpret subsequent cosmological processes accurately. Consequently, studies of primordial neutrino asymmetries often rely on approximations, such as the use of thermal Fermi-Dirac distributions for (anti)neutrinos, without rigorous verification.
@@
In this work, we address these ambiguities by numerically solving the neutrino decoupling process and analyzing the resulting (anti)neutrino density matrices in the context of BBN, CMB, and LSS. We rederive the quantum kinetic equations (QKEs) for neutrinos and antineutrinos using the Closed Time Path (CTP) formalism for two-point functions. Assuming complete flavor equilibration before neutrino decoupling, where $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$, we numerically solve the QKEs in the early universe using an extended version of the public code FortEPiaNO Gariazzo:2019gyi ; Bennett:2020zkv . Our calculations include the effects of neutrino oscillations, matter interactions, FTQED corrections up to $\mathcal{O}(3)$, and comprehensive neutrino-electron and neutrino-neutrino collision terms. The final $N_{\rm eff}$ is then computed based on the resulting density matrices and can be expressed as $N_{\rm eff}\simeq N_{\rm eff}^{\rm SM}+3\left(\frac{30}{7\pi^{2}}\xi_{\nu}^{2}+\frac{15}{7\pi^{4}}\xi_{\nu}^{4}\right)+0.0102\xi_{\nu}^{2}$, where $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$. The second and third terms represent corrections from nonzero $\xi_{\nu}$ in the instantaneous decoupling limit and non-instantaneous effects, respectively. We also confirm that spectral distortions of neutrinos and antineutrinos with nonzero $\xi_{\nu}$ differ due to distinct Pauli-blocking effects, in agreement with preliminary studies Grohs:2016cuu .
@@
Using the neutrino and antineutrino density matrices, we further investigate the effects of primordial neutrino asymmetries on BBN, CMB, and LSS. For BBN, we introduce a novel correction to the weak interaction rates arising from distinct neutrino and antineutrino spectral distortions. We calculate the light element abundances using the public BBN code PRIMAT Pitrou:2018cgg , incorporating additional corrections to weak rates. According to the EMPRESS measurement of the helium abundance $Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033}$, our results suggest a positive primordial neutrino asymmetry in the range $0.032\leq\xi_{\nu}\leq 0.052$, assuming a fixed baryon abundance $\omega_{b}=0.024$. Our analysis also highlights the importance of spectral distortions for precise BBN predictions, particularly for Helium-4 abundance. Compared to results in the instantaneous limit, the traditional treatment using Fermi-Dirac distributions predicts a negative relative difference for positive primordial neutrino asymmetry, whereas our precise treatment with spectral distortions shows a positive relative difference. Additionally, we provide semi-analytic expressions for these abundances as functions of $\xi_{\nu}$, which will be valuable for future studies. For CMB and LSS, our study reveals that primordial neutrino asymmetries influence the amplitudes of acoustic oscillations in both the CMB and baryon acoustic oscillations (BAO), suggesting that BAO observations can serve as a tool for constraining these asymmetries.
@@
To further constrain primordial asymmetries, we perform a joint MCMC analysis including the neutrino degeneracy parameter $\xi_{\nu}$ and the total neutrino mass $\sum m_{\nu}$, using data from EMPRESS BBN, Planck CMB, and BOSS BAO observations. Combining the EMPRESS BBN and Planck CMB datasets, we constrain the primordial neutrino asymmetry to $\xi_{\nu}=0.028\pm 0.012$, a reduction of $\sim 18\%$ in the central value and a slight reduction in the uncertainty compared to previous results of $\xi_{\nu}=0.034\pm 0.014$ Escudero:2022okz . Incorporating BOSS BAO observations further refines the constraint to $\xi_{\nu}=0.024\pm 0.012$, yielding a $\sim 29\%$ reduction in the central value and a $2.0\sigma$ significance for a positive neutrino asymmetry. We also discuss the interplay between $\sum m_{\nu}$ and $\xi_{\nu}$ when both are treated as free parameters, and how these constraints translate into bounds on UV model parameters, using a specific model where large neutrino asymmetries are generated by Q-ball decay Kawasaki:2002hq ; Gelmini:2020ekg .
@@
Our work establishes a framework for exploring the implications of physics beyond the Standard Model and $\Lambda$CDM for neutrino decoupling and related cosmological processes, providing corresponding constraints from current and future BBN, CMB, and LSS observations. It paves the way for more precise studies of primordial neutrino asymmetries and other new physics scenarios, such as synchronous neutrino oscillations Froustey:2021azz ; Froustey:2024mgf and non-standard neutrino interactions Du:2021idh ; Du:2023upj , driven by increasingly precise cosmological observations.